OpenStudy (anonymous):

High lvl math: Find the parabola y=ax^2+bx+c which intersects the x-axis at the points x=3 and x=-5, and the y-axis at the point y=30

OpenStudy (amistre64):

determine the general solution with those roots: (x-r1)(x-r2) = x^2-(r1+r2)x + r1r2

OpenStudy (amistre64):

when x=0, we have y = r1r2 which needs to be scaled to equal 30

OpenStudy (anonymous):

You have c=30, now use your x values to get simultaneous equations in a and b.

OpenStudy (amistre64):

x^2-(r1+r2)x + r1r2 ; using r1= 3 and r2 = -5 x^2 -(3--5)x + (3*-5) x^2 -8x -15 ; at x=0; y=-15 ; -15s = 30 when s=-2 y = -2(x^2 -8x -15)

OpenStudy (amistre64):

hmm, im off a bit

OpenStudy (amistre64):

i got me roots off somehow :) let me dbl chk

OpenStudy (anonymous):

I get (-2x-10)(x-3) which can be arranged a bit, obviously.

OpenStudy (amistre64):

-2(x-3)(x+5)

OpenStudy (anonymous):

I prefer your method myself but the simul eqs might be a bit easier to understand.

OpenStudy (anonymous):

thanks for help :)

OpenStudy (amistre64):

x^2 -(3--5)x + (3*-5) ^ thats spose to be a +- :)

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

estudier can you show me your work if you have time because I got something wrong